LBT and LUCI Technical Observing, prep for Partner Science:
UT 2011 Jun 24

Observer: DThompson (KBoutsias and RSpeziali also present)
Telescope Operator: GBechetti
Instrument Support: JMorris
Telescope Support: JLittle
SW support: NCushing

Summary

LBC

LUCI

Preparations

Details

LBC

01:52 Shutdown UI and reboot CMU
02:03 Warmup for a couple minutes and then turn on
02:08 10Bias_Bino (not dark in dome)
02:16 2Darks_Bino (not dark in dome, moved telescope during)
02:50 SkyFlatTest_VV made with mkskyflat - saturated
02:55 SkyFlatTest_gr - still saturated
03:02 SkyFlat_gr 0.5, 1.0, 2.0 scalings
  • B 030104 24k @ start
  • R 030101 15k @ start
03:14 Sky flat scaling... (OBs need to be made generic!). Taken too late?
  • Blue_SkyFlatScalings.ob
  • Blue_SkyFlatScalings.ob with scale=5.0
  • Blue_SkyFlatScalings2.ob with scale=5.0
  • Red_SkyFlatScalings.ob with scale=10.0
  • Red_SkyFlatScalings2.ob with scale=10.0
  • Note: using a scale factor consistently gives a popup error box stating "OB upload error: 4", ignored.
  • Note2: sources are significantly out of focus.

03:43 Pointing check on ACT0311 (EL~83)
  • lbcrangebal hit limits! Cleared all active optics, instrument offsets, etc. Left in -1mm Z on both
  • lbcrangebal 34942 34940, worked!

03:55 DOFPIA, /first @ ACT0311 (may not be enough stars) - failed, no usable pupils

04:02 DOFPIA, first @ ACT0298 Red failed to converge (rejected most pupils)

04:15 DOFPIA, @ ACT0247 working, exited with error (command rejected since another is running).

04:28 DOFPIA at Pal14 field...again an error (command rejected since another is running). Manual recovery is to dismiss the error message and trigger the OB by hand. Dofpia is just waiting for the images to appear in /newdata. Red is taking forever to converge! Both sides finally converged after ~12 cycles. All together it took 67 minutes to converge.

05:02 Dither test on INAF field. Dithers are working. Guiding appears to be oscillating quite a bit. Thjere was a telescope jump in the RED data (lbcr.20110624.052240). The blue image (052242) also looks affects but appears to have guided most of it out successfully (round cores with wings in the direction of the jumps).

05:48 Measure focal plane tilt. EL~70
  • dofpia to recollimate (3 iterations)
  • RB_rVsuperfoc.ob
  • RB_rVsuperfoc180.ob
  • dofpia to recollimate (3 iterations)

06:02 Repeat dither test on INAF field. Guiding still looks like it is oscillating, like the red and blue channels are pushing each other around. Need to investigate the binocular LBC guiding algorithm. The only way for the two to affect each other is if some/all of the guiding offsets are applied to the mount. IT# (critical) Of the six 600s exposures, five are unusable.
  • BLUE RED
  • 050034 050031
  • 051148 051149
  • 052242 052240
  • 060141 060135
  • 061228 061230
  • 062322 062324

Correcting pointing/co-pointing on ACT0247 Ran unguided exposures fo 60 and 180s - perfectly round. There is a problem with guiding, either LBC is sending bad data or the telescope is applying it incorrectly.

06:57 Switch to LUCI.

LUCI

System time is OK
Mask tables are OK
02:14 After restart of SW, MOS Sequence Server did not restart. Started by hand.
02:33 MOS Sequence Server started by hand. Initialize on UI successful. 02:42 Sieve mask image in N3.75 OK; poor on bottom third of N1.8 07:19 Calibration unit & halo lamps verified
07:21 LUCI1 Daily functional checkout: passed
    • O2DCR 2.0s (usually around 11.6 e-) 0009&10 RN=11.78 (normal)

07:25 Correct initial collimation/pointing (Geno)

Re-test updatePointingReference functionality (high priority)
  • BS9183 at PA=30 preset before each block test below...
  • I] REL/RADEC offset followed by RADEC absorb. 0014-0025
  • II] REL/DETXY offset (like shift image) followed by RADEC absorb. 0026-0038
  • III] REL/RADEC & REL-DETXY offsets followed by RADEC absorb. 0039-0051
  • IV] REL/RADEC offset followed by DETXY absorb. 0052-0064
  • V] REL/DETXY offset followed by DETXY absorb. 0065-0076
  • VI] REL/RADEC & REL-DETXY offsets followed by DETXY absorb. 0077-0090
  • VII] REL/RADEC offset followed by DETXY & RADEC absorbs. 0091-0102
  • VIII] REL/DETXY offset followed by DETXY & RADEC absorbs. 0103-0114
  • IX] REL/RADEC & REL_DETXY offsets followed by DETXY & RADEC absorbs. 0115-0128

Take set of linearization test data (photometric weather only!)
  • Aborted before the 32s exposures because the 27s are saturated in the cores
  • LUCI images: 0129-0169

09:42 MOS acquisition tests. Poorly prepared...only three alignment sources, two of which are not visible on 30s K images!

10:15 MOS error - FPU clamp x failed to reach negative limit when clamping mask #17 into the FPU. System is in an almost safe state. Peter walking me through recovery.

End of night or tomorrow afternoon:
  • Take 300s dark w/z filter to monitor "blob"
  • Run LUCI1 flexure tests, see: ~/LBTO/Flexures/RotatorScript.txt (closed dome)
  • Test sound playing through IIF (write iraf task to do this)

-- DavidThompson - 24 Jun 2011
Topic revision: r3 - 24 Jun 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback